OpenStudy (anonymous):

anybody can remember how to do this? find the value of the unknown variables. x+y=24 y+z=26 x+z=28 i badly need. .

OpenStudy (anonymous):

@lgbasallote

OpenStudy (lgbasallote):

x = 24 - y 24 - y + z = 28 -y + z = 4 use that with the second equation y + z = 26 -y + z = 4 2z = 30 z = 15 got that?

OpenStudy (anonymous):

does it has the same process in finding the other variables?